OpenStudy (madgirlwithabluebox):

Help! Will Fan and give medal! Find any points of discontinuity for the rational function. y= (x+6)(x+2)(x+8)/(x+9)(x+7)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

set the denominator equal to zero and solve for \(x\) i.e. solve \[(x+9)(x+7)=0\] (pretty much in your head)

OpenStudy (madgirlwithabluebox):

Could you give me a hint as to what to plug in for x? its early and i cant think right :(

OpenStudy (misty1212):

\[x+9=0\]or \[x+7=0\] is the next step

OpenStudy (madgirlwithabluebox):

ohkay i think im over thinking this... would it just x= -9,-7 ?

OpenStudy (misty1212):

yup

OpenStudy (madgirlwithabluebox):

so what do i do now?

OpenStudy (misty1212):

go have a soda, you are done

OpenStudy (madgirlwithabluebox):

so x= -9 x=-7are the points of discontinuity?